Hi,
I do not see any dection for the problem I am asking for a solution for. According to my book this is: Complete the ordered pairs so that each is a solution for the given equation. 3x + 4y = 12 (0, ),( ,3/4),( ,0),(8/3, ).
I am not sure if I wrote the fractions correctly 3/4 and 8/3.


Thank you for your help


Answer by algebrapro18(249)

well all you have to do is plug in the number for the correct varialbe and solve.
pluging 0 in for x
3(0)+4y=12

simplifying
4y=12

dividing both sides by 4
y=3

(0,3)



plinging 3/4 in for y
3x+4(3/4)=12

simplifying
3x+3=12

subtracting 3 from both sides
3x=9

dividing both sides by 3
x=3

(3,3/4)


Pluging in 0 for y you get
3x+4(0)=12

simplifying
3x=12

dividing both sides by 3
x=4

(4,0)


pluging in 8/3 for x you get
3(8/3)+4y=12

simplifying
8+4y=12

subtracting 8 from both sides
4y=4

dividing both sides by 4
y=1

(8/3,1)